OpenStudy (anonymous):

What Percent of 48 is 20?

OpenStudy (alexandervonhumboldt2):

to find this \[48*\frac{ x }{ 100 }=20\]

OpenStudy (anonymous):

im lost bro

OpenStudy (anonymous):

i have 3 hours finding this answer man =(

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

just solve for X

OpenStudy (anonymous):

please alex i don't get it please brother private or something man

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

@Mehek14

OpenStudy (anonymous):

@Willie579

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

akexa i have all day trying to solve this

OpenStudy (alexandervonhumboldt2):

well , to find the number percent to need to multply the number by number of percents and get the percent. 48*x/100=20 x=41.6 %

OpenStudy (anonymous):

i went to that site 20 times already

OpenStudy (willie579):

20 is 41.67% of 48.

OpenStudy (anonymous):

that wrong

OpenStudy (anonymous):

no

alones (alones):

it should have been 48*x/100=20 x=41.67 and then 100-41.67=58.3 so that is your answer

OpenStudy (anonymous):

wrong

OpenStudy (anonymous):

nvm i skip the question ima not getting credit for it fk it